OpenStudy (anonymous):

Can someone please help me with graphing linear equations? I just don't understand it, I always get the points mixed up, and put them in the wrong spots. I don't really have a clue of what I'm doing.

OpenStudy (anonymous):

1. start with x = 0 then x = 1, x =2, etc. 2. plug it into the equation and you should get y 3. now you have your x coordinate (from step 1) and your y coordinate (from step 2).

OpenStudy (anonymous):

Sorry, I'm still confused... Maybe I should draw out an example problem.

OpenStudy (anonymous):

ok, hold on

OpenStudy (anonymous):

for example you get the equation y = 2x + 3 you start getting values for x. lets start with x = 0. you plug it in 2(0) + 3 = 3. 3 is your value for y so your first coordinate is (0, 3) next we do x = 1. plug it in 2(1) + 3 = 5. now now you have another point. (1, 5) x = 2 -> 2(2) + 3 = 7. 3rd coordinate is (2, 7) you keep increasing the values of x and keep finding the points and then it'll form a line

OpenStudy (anonymous):

I'm sort of getting it. Here's a problem though.. just a sec

OpenStudy (anonymous):

y= x-1 y+-x+3 Graph y=x-1 by plotting the y intercept, a point based on a slope of 1.

OpenStudy (anonymous):

wish I could have some electronic graphing paper.

OpenStudy (anonymous):

same concept, x = 0 plug it in to y = x -1 0 - 1 = -1 so x = 0, y = -1 for x = 1, 1-1 = 0 so x = 1, y = 0 x = 2, 2-1 = 1 so x = 2 and y = 1

OpenStudy (anonymous):

Thank you so much! This makes so much sense now! I just had to think about it.